Q.1 A body is moved along a str. line by a machine delivering a constant power .  The distance travelled by the body in time t is proportional to
  1. t ^3/4
  2. t^3/2
  3. t^1/4
  4. t^1/2                                                   Q2. The potential energy of a 1 kg particle free to move along the x – axis is given by ( X^4/4 – X^2/2) ,then the maximuim speed in m/s is                    a 3/ squareroot 2                                  b. 2^ ½                                             c. 2            

Grade:11

1 Answers

Nirmal Singh.
9 years ago
Dear Students Please post question one by one.

Ques.1 – -

Power Deliver is constant so P = constat

From work energgy theoru, we know that W = change in kinetic Energy = 0.5 m v2

So Power P = W / t

P = 0.5 m v2 / t

v= sqrt (2 Pt / m)

v = dx / dt =d / dt [sqrt (2 Pt / m)]

integral of dx = intergral ofsqrt (2 Pt / m) dt

so x = 2 /3 sqrt (2P/m) t3/2

Hence x is directly proprtaional to t3/2

Option (2) is correct

Regards,
Nirmal Singh
Askiitians Faculty